OpenStudy (anonymous):

How can x + 2 = -3x + 11 be set up as a system of equations?

OpenStudy (anonymous):

are you trying to find x?

OpenStudy (anonymous):

Nope im just trying to set it up as a equation,

OpenStudy (anonymous):

i think it's y = x + 2 y = -3x + 11

OpenStudy (anonymous):

@Awesome781 @beccaboo333 @bibby and @Compassionate can you guys check this out because i am not sure and dont want to lead them the wrong way

OpenStudy (bibby):

If you try substitution you van replace the y in the first equation with -3x + 11 and solve for x I believe

OpenStudy (anonymous):

im not trying to solve for X i just have to set it up as a system of equations, a) y = x + 13 y = -3x + 11 b) y = x + 2 y = -3x + 11 c) y = x + 2 y = -2x + 11 d) y = 2 y = -3x + 11

OpenStudy (anonymous):

I believe its b) y = x + 2 y = -3x + 11

OpenStudy (bibby):

That seems right

OpenStudy (anonymous):

it looks correct to me

OpenStudy (anonymous):

alright thanks......not sure who the best answer go's to o.o

OpenStudy (anonymous):

thank you @bibby for helping

OpenStudy (anonymous):

we will make it fair and we all get medals. I gave you a medal @Futuristicmario because you tried to answer the question and thats what we like

OpenStudy (anonymous):

lol thanks.